# Look Before you Leap: Estimating LLM Benchmark Scores from Descriptions

### v2

@@ 5 Experiments
We compare against three families of baselines. First, we report the overall test-split mean across all metrics (Section 5.1) and the per-metric test-split means (Section 5.2). Because no training or validation set is available, the test-set average serves as a strong baseline for comparison. Second, we evaluate text-encoder–based regressors. We embed each description and train classical models—k-nearest neighbors (guo2003knn) and XGBoost (chen2016xgboost)—on these embeddings. With no dedicated training set, we use k-fold cross-validation (fushiki2011estimation): fit on k-1 folds, evaluate on the held-out fold, and average over folds. We set k = 10 and used E5-Mistral-7B (wang2023improving), which is a strong dense embedding encoder baseline. Hyperparameters were tuned on a held-out validation split from the 2023–2024 version of Precog.
@@
We evaluate proprietary and open-source LLMs in both reasoning and non-reasoning modes to disentangle chain-of-thought effects from base generation quality. On the proprietary side, we use GPT-5 (openai2025gpt5) in its reasoning configuration; on the open-source side, we include reasoning-enabled Qwen3-32B (yang2025qwen3) and non-reasoning Qwen2.5-32B/72B (team2024qwen2). For reasoning-capable models, we also run a no-reasoning variant (GPT-5 with minimal reasoning; Qwen3 with no-thinking) to quantify the contribution of explicit reasoning. We match prompts and decoding settings across modes so that performance differences can be attributed to reasoning. In all settings, the model produces a brief rationale and encloses its final estimate in boxed{}, which we parse in post-processing. GPT-5 uses default decoding hyperparameters, and Qwen3 uses the original authors’ recommended settings.
@@ 5.1 Evaluation Results
Table 3: Prediction performance on Precog for 2023–2024 (767) and 2025 (1,523 instances). indicates significance at $p<0.05$ (one-sided binomial sign test) for the corresponding column.
We summarize the main validation results in Table 3. With the exception of GPT-5 (openai2025gpt5), no baseline meaningfully outperforms the test-set global mean. GPT-5 achieves the lowest overall MAE (14.6), but still leaves substantial headroom. Performance is similar on the 2023–2024 and leakage-free 2025 splits, suggesting that prediction accuracy is not driven by training-data memorization. Moreover, reasoning ablations for GPT-5 and Qwen3 show that enabling reasoning mode improves performance prediction. Classical text-encoder baselines—kNN and XGBoost with k-fold CV—also fail to beat the global mean by a meaningful margin. Among Qwen variants (team2024qwen2; yang2025qwen3), Qwen3-32B in reasoning mode performs best, and is comparable to Qwen2.5-72B despite using roughly half as many parameters.
5.2 Prediction Heterogeneity by Task, Metric, Target, and Setup
We evaluate GPT-5 across seven metric groups by reporting per-metric MAE relative to a test-calibrated per-metric mean baseline, which is stronger than a single global mean. As shown in Figure 3, GPT-5 typically outperforms this baseline on both the 2023–2024 and 2025 splits. The Accuracy subset—our largest slice—closely tracks the overall trend. In contrast, for Rouge and Precision, whose gold scores have low variance, LLMs sometimes fall below the baseline, underscoring the difficulty of forecasting such fine-grained metrics.
Figure 3: Per-metric MAE (↓) by model, grouped by evaluation metric and computed separately on the 2023–2024 and 2025 data subsets. The red and orange bars show the per-metric mean, labeled with its absolute MAE; the other bars show deviations from that mean within the corresponding year-split subset.
@@ 5.2 Prediction Heterogeneity by Task, Metric, Target, and Setup
(a) By benchmark category.
@@
(b) Peer-reviewed vs. non–peer-reviewed.
@@
(d) Established datasets vs. New datasets.
We analyze heterogeneity on the Precog 2025 snapshot along benchmark category, publication status, target model, and established datasets vs. new datasets in Table 4. We define established datasets records as instances where the result reporting paper differs from the dataset paper, and new datasets records as instances where they are the same (the result paper introduced the dataset). Across categories, both GPT-5 and Qwen3-32B show similar patterns: multilingual benchmarks are easiest to forecast (lowest MAE, highest $r$), while math and instruction-following are harder, likely reflecting more diverse task formulations and evaluation setups. Performance is also very similar on peer-reviewed and non–peer-reviewed benchmarks, with slightly better MAE on non–peer-reviewed entries, suggesting that venue and “maturity” have limited impact.
Conditioning on the target model, both predictors perform better on GPT-4 than GPT-4o, indicating that newer or more variable targets are modestly harder to forecast, plausibly because GPT-4 results were more exposed during LLM training. Finally, MAE and correlation are nearly identical for established datasets vs. new datasets records, implying that the predictors transfer well from established datasets to newly introduced, previously unseen datasets.
@@ 5.3 Human prediction results
Model MAE $\mathrm{r}$
@@
Table 5: Human Prediction Results on Accuracy subset ($n=140$).
In the previous sections, we demonstrated that LLMs can generate accurate predictions from dataset descriptions. A natural next question is how these predictions compare to those of human experts, who are known to struggle with probability judgments and decisions under uncertainty (kahneman1972subjective). To explore this, we asked six graduate students and one professor of computer science, all specializing in NLP and machine learning, to each predict 20 accuracies based solely on the same text descriptions provided to models. On this 140-item subset, human predictions correlated moderately with ground truth ($r=0.403$) and outperformed the mean baseline in terms of MAE (19.6 vs. 21.7). Nevertheless, GPT-5 achieved substantially stronger results, indicating that while humans can do the task reasonably well, advanced reasoning models achieve markedly lower MAE. Although the study has a small sample size, the results primarily serve to illustrate the inherent difficulty of the task. While human forecasters might improve with additional training (tetlock2014forecasting), the fact that LLMs already outperform researchers on this performance prediction task suggests that benchmark prediction models have the potential to provide meaningful support to practitioners when planning experiments.
@@ 6 Confidence Calibration
Some experimental outcomes may be inherently difficult to predict, raising the question of whether models can reliably estimate their own prediction accuracy. To examine this, we elicit a scalar, self-reported confidence for each prediction (following yoon2025reasoning) and evaluate calibration by filtering to items with confidence $\geq\tau$. For each $\tau$, we report the MAE on the retained subset together with its coverage. This gives a simple view of the error–coverage trade-off. We report confidence curves for the accuracy subset, which dominates Precog.
@@ 7 Retrieval Analysis
We assess how retrieval contributes to performance prediction using tool-call counts and query token diversity (tokenized with GPT-4o (hurst2024gpt)). All models share a custom retriever over the arXiv API, [^2] https://info.arxiv.org/help/api/index.html which returns canonical arXiv PDF links. To avoid leakage, we exclude each record’s source papers by identifier during retrieval. We initially tried GPT-5’s built-in web search—which is more token-efficient but corpus-uncontrolled—and found that it frequently surfaced source papers: in an audit, roughly 70% of evaluation source papers became identifiable, so we instead use the custom retriever for GPT-5 despite its higher cost. We report results with retrieval enabled and disabled to measure gains from literature grounding beyond the model’s prior. Each query returns at most two documents, and the agent may issue up to four retrievals under ReAct-style prompting (yao2023react).
@@ 8 Streaming Prediction
Figure 5: Streaming prediction error boxplots of GPT-5+Web-Search.
@@
We collect candidates from arXiv’s official interfaces (API/RSS) for cs.CL, cs.LG, and cs.AI, and fetch PDFs via the recommended hosts, respecting robots.txt and published rate limits; we also notified arXiv staff in advance (see Appendix G).
A daily pipeline then (i) identifies papers of interest, (ii) extracts table records, (iii) builds redacted descriptions, and (iv) issues predictions—end-to-end before third-party search engines index the papers. In this setting, we restrict to records whose metrics map to our seven metric groups and expand the model axis to the GPT family. We permit built-in web-search here, as leakage from prior literature is negligible for freshly released papers. We collected 52 papers containing 79 previously unseen experimental records and datasets between 2025-09-05 and 2025-09-23. During search, GPT’s web tool may surface the target paper. We match arXiv IDs and verify content to detect any such access, and we exclude those predictions to prevent leakage.
Model MAE $\mathrm{r}$
@@
Table 7: Streaming environment prediction results on 79 samples.
@@ Results
Table 7 summarizes streaming performance for GPT-5 (+Search). On newly released datasets and unseen experiment–dataset combinations, the model achieves accuracy comparable to its performance on Precog (MAE =14.0). The box plot of absolute errors (Figure 5) compares all predictions with the high-confidence subset (>>50% confidence, 23 samples), showing a visibly tighter error distribution for the latter (avg. 9.3). Finally, in a self-referential test that predicted the performance of the task described in this paper (not yet on arXiv), GPT-5+Search predicted its own performance at MAE 16.0 versus the realized 14.7 from Table 3. These results indicate that forecasting performance based solely on text-only descriptions could be a practical and achievable approach in real-world applications.
9 Conclusion & Discussion
Conclusion.
We recast evaluation as forecasting and introduce Precog, a leakage-mitigating corpus of redacted, schema-aligned description–performance pairs. GPT-5 produces moderately accurate forecasts with calibrated uncertainty, and remains effective in a streaming setting where results are predicted for newly released papers prior to broad indexing. These findings indicate that anticipatory evaluation is feasible and practically useful for estimating task difficulty, prioritizing experiments, and reducing compute.
Discussion: scope and limitations.
[…]
Discussion: cost and latency.
Using GPT-5 (medium reasoning, no search) via the Batch API costs about $0.01$ per query and averages $\sim$132 seconds per prediction with high parallelism. For open-source baselines (e.g., Qwen), we run single-GPU inference on an A40 with a maximum generation length of 32,76832{,}768 tokens; a single prediction averages $\sim$48 seconds. While not directly comparable to end-to-end benchmark runs, description-based forecasting is markedly more cost- and time-efficient, making it well suited for early screening and study design.
Discussion: limits of current open-source models.
Despite progress, open-source baselines lag behind GPT-5 in both accuracy and calibration. They show brittle tool use (inconsistent gains from retrieval), weaker representations of fine-grained cues in long descriptions, and higher variance across metric families (e.g., small-variance metrics like BLEU/ROUGE). Improved reasoning and uncertainty modeling are needed for reliable deployment.
Outlook: toward specialized forecasting models.
To accelerate scientific discovery, we advocate purpose-built regressors that leverage LLM reasoning. In particular, reasoning-aware regression trains a regression model on top of reasoning LLMs, while better use of retrieval tools focuses on learned query planning and selective evidence use. An important future direction is to make these forecasting systems robust for new models with very few or no prior evaluations, for example by incorporating stronger priors, actively collecting a small set of seed evaluations, or using richer hybrid text–metadata representations. Developing these components on top of Precog and the streaming benchmark should yield more accurate, trustworthy, and cost-efficient forecasting systems.
10 Ethics Statement
We rely exclusively on publicly available research artifacts (e.g., arXiv PDFs and published results); no human subjects or personally identifiable data are involved, and all usage follows licenses and provider terms. To reduce leakage, we redact dataset identities and block retrieval of each record’s table-source and dataset-source papers; because residual cues may persist, we also evaluate a time-gated, zero-leakage setting on genuinely unseen datasets. Focusing on arXiv means some records lack full experimental details, which can introduce label noise; to aid auditing, we release venue metadata per record and document known issues. Our strongest results use large proprietary models, raising equity and environmental concerns; we therefore report compute, provide open baselines, and encourage reproduction with smaller models. Finally, the corpus is predominantly English, limiting generality; expanding multilingual coverage is an important direction for future work.
11 Reproducibility Statement
We will release all artifacts needed for exact or statistically equivalent reproduction: the full Precog curation stack (scripts, leakage filters, redaction utilities), the forecasting pipelines (retrieval, prompting, decoding, scoring), and the evaluation scripts. Our retriever queries the arXiv API, returning at most two PDFs per query and up to four queries per instance (ReAct-style); we will also release the exact query templates and our custom arXiv wrapper. To model unseen datasets, we enforce a zero-leakage protocol by time-gating retrieval to snapshots before each dataset’s index date, and we publish per-record snapshot dates plus a script. All prompts (system/instruction/user and ReAct) are released verbatim; decoding parameters are fixed and logged, defaulting to greedy where configurable. For GPT-5, decoding settings are provider-controlled; we use defaults and record access dates. Baselines (E5-Mistral + kNN/XGBoost; Qwen3 variants) are used off-the-shelf; we provide exact checkpoints or API identifiers and access dates, and no baseline is finetuned. Upon acceptance, we will release code, data, prompts, configs, and logs under <MIT LICENSE>; non-redistributable third-party assets are referenced with exact versions and retrieval instructions.
@@ Appendix A Details of Data Curation
For curation we used GPT-5-mini (openai2025gpt5). The model extracted schema-aligned descriptions from source papers and handled dataset-source retrieval annotations. These choices were based on a pilot run and seed-result quality.
@@
All records involve GPT-4 (achiam2023gpt) or GPT-4o (hurst2024gpt), fixing the model/method axis to focus on dataset/task variation. We excluded multimodal domains, correlation-only metrics (not on a 0–100 scale), and overall benchmark aggregates (e.g., MMLU) via heuristic filters.
@@ Appendix C Supplementary Experimental Results
We ran additional experiments on the Precog-2025 dataset using more recent embedding models that are competitive or state-of-the-art on the MTEB leaderboard, including Qwen3-8B-Embed (zhang2025qwen3) and KaLM-Embedding-Gemma3-12B-2511 (hu2025kalmembedding), combined with both KNN and XGBoost regressors. The results are summarized in Table 8. As the table shows, newer and stronger embedding models yield modest improvements in MAE and correlation, but the gains are relatively small.

### v3

@@ 5 Experiments
We compare against three families of baselines. We report the overall test-split mean across all metrics (Section 5.1). Because no training or validation set is available, the test-set average serves as a strong baseline for comparison. Second, we evaluate text-encoder–based regressors. We embed each description and train classical models—k-nearest neighbors (Guo et al., 2003) and XGBoost (Chen and Guestrin, 2016)—on these embeddings. With no dedicated training set, we use k-fold cross-validation (Fushiki, 2011): fit on k-1 folds, evaluate on the held-out fold, and average over folds. We set k = 10 and used E5-Mistral-7B (Wang et al., 2023), which is a strong dense embedding encoder baseline. Hyperparameters were tuned on a held-out validation split from the 2023–2024 version of Precog.
@@
We evaluate proprietary and open-source LLMs in both reasoning and non-reasoning modes to disentangle chain-of-thought effects from base generation quality. On the proprietary side, we use GPT-5 (OpenAI, 2025) in its reasoning configuration; on the open-source side, we include reasoning-enabled Qwen3-32B (Yang et al., 2025) and non-reasoning Qwen2.5-32B/72B (Team, 2024). For reasoning-capable models, we also run a no-reasoning variant (GPT-5 with minimal reasoning; Qwen3 with no-thinking) to quantify the contribution of explicit reasoning. In all settings, the model produces a brief rationale and encloses its final estimate in boxed{}. GPT-5 uses default decoding hyperparameters, and Qwen3 uses the original authors’ recommended settings.
Figure 3: Per-metric box plots of absolute errors in GPT-5 predictions on the 2025 dataset.
@@ 5.1 Evaluation Results
Table 3: Prediction performance on Precog for 2023–2024 (767) and 2025 (1,523 instances). indicates significance at $p<0.05$ (one-sided binomial sign test (Hollander and Wolfe, 1999)) for the corresponding column.
@@
(a) By benchmark category.
@@
(b) Peer-reviewed vs. non–peer-reviewed.
@@
(c) By target model (GPT-4 vs. GPT-4o).
@@
We summarize the main validation results in Table 3. With the exception of GPT-5 (OpenAI, 2025), no baseline meaningfully outperforms the test-set global mean. GPT-5 achieves the lowest overall MAE (14.6), but still leaves substantial headroom. Performance is similar on the 2023–2024 and leakage-free 2025 splits, suggesting that prediction accuracy is not driven by training-data memorization. Moreover, reasoning ablations for GPT-5 and Qwen3 show that enabling reasoning mode improves performance prediction. Classical text-encoder baselines (kNN and XGBoost) also fail to significantly beat the global mean. Among Qwen variants (Team, 2024; Yang et al., 2025), Qwen3-32B in reasoning mode performs best, and is comparable to Qwen2.5-72B despite using roughly half as many parameters.
5.2 Prediction Heterogeneity by Different Subset
We analyze per-metric prediction behavior by plotting metric-wise absolute error box plots on Precog 2025. As shown in Figure 3, GPT-5 exhibits moderate absolute errors in accuracy and exact match metric. However, for recall and precision, the model shows comparatively larger errors, with both the median and mean higher than the other metrics. This may be because relatively few training records include these metrics compared to accuracy, which could limit the model’s ability to predict them accurately. Also, errors on BLEU and ROUGE tasks are smaller, likely due to the characteristics of these metrics; within this group, ROUGE errors are higher than BLEU, possibly because BLEU is primarily used for translation evaluation, whereas ROUGE is applied across a more diverse set of tasks.
We also look into heterogeneity on the Precog 2025 snapshot along benchmark category, publication status, target model, and established datasets vs. new datasets in Table 4(d). We define established datasets records as instances where the result reporting paper differs from the dataset paper, and new datasets records as instances where they are the same.
Across categories, both GPT-5 and Qwen3-32B show similar patterns: multilingual benchmarks are easiest to forecast, while math and instruction-following are harder, likely reflecting more diverse task formulations and evaluation setups. Performance is also very similar on peer-reviewed and non–peer-reviewed benchmarks, suggesting that venue and “maturity” have limited impact. Conditioning on the target model, both predictors perform better on GPT-4 than GPT-4o, indicating that newer targets are modestly harder to forecast, plausibly because GPT-4 results were more exposed during training. Finally, MAE and correlation are nearly identical for established datasets vs. new datasets records, implying that the predictors transfer well from established datasets to unseen datasets.
@@ 5.3 Human prediction results
We demonstrated that LLMs can generate accurate predictions from dataset descriptions. A natural next question is how these predictions compare to those of human experts, who are known to struggle with probability judgments and decisions under uncertainty (Kahneman and Tversky, 1972). To explore this, we asked six graduate students and one professor of computer science, all specializing in NLP and machine learning, to each predict 20 accuracies based solely on the same text descriptions provided to models. The results are shown in Table 5 where human predictions correlated moderately with ground truth ($r=0.403$) and outperformed the mean baseline in terms of MAE (19.6 vs. 21.7). Nevertheless, GPT-5 achieves substantially stronger results. This indicates that while human experts perform reasonably well, the reasoning model attains a markedly lower MAE.
Table 5: Human Prediction Results on Accuracy subset ($n=140$).
Model MAE $\mathbf{r}$
@@
Although the study has a small sample size, the results primarily serve to illustrate the inherent difficulty of the task. While human forecasters might improve with additional training (Tetlock et al., 2014), the fact that LLMs already outperform researchers on this performance prediction task suggests that models have the potential to provide meaningful support to practitioners when planning experiments.
@@ 6 Confidence Calibration
Some experimental outcomes may be inherently difficult to predict, raising the question of whether models can reliably estimate their own prediction accuracy. To examine this, we elicit a scalar, self-reported confidence for each prediction (following Yoon et al. (2025)) and evaluate calibration by filtering to items with confidence $\geq\tau$. For each $\tau$, we report the MAE on the retained subset together with its coverage, which gives a view of the error–coverage trade-off. We report confidence curves for the accuracy subset, which dominates Precog.
@@ 7 Retrieval Analysis
We assess how retrieval contributes to performance prediction using tool-call counts and query token diversity (tokenized with GPT-4o (Hurst et al., 2024)). All models share a custom retriever over the arXiv API, [^2] https://info.arxiv.org/help/api/index.html which returns canonical arXiv PDF links. To avoid leakage, we exclude each record’s source papers by identifier during retrieval. We initially tried GPT-5’s built-in web search—which is more token-efficient but corpus-uncontrolled—and found that it frequently surfaced source papers: in an audit, roughly 70% of evaluation source papers became identifiable, so we instead use the custom retriever for GPT-5 despite its higher cost. We report results with retrieval enabled and disabled to measure gains from literature grounding beyond the model’s prior. Each query returns at most two documents, and the agent may issue up to four retrievals under ReAct-style prompting (Yao et al., 2023).
@@ 8 Streaming Prediction
Figure 5: Streaming prediction error boxplots of GPT-5+Web-Search. Confident denotes the prediction subset with over 50% confidence (Yoon et al., 2025).
@@
We collect candidates from arXiv’s official interfaces (API/RSS) for cs.CL, cs.LG, and cs.AI, and fetch PDFs via the recommended hosts, respecting robots.txt and published rate limits (see Appendix G).
A daily pipeline then (i) identifies papers of interest, (ii) extracts table records, (iii) builds descriptions, and (iv) issues predictions—end-to-end before third-party search engines index the papers. In this setting, we restrict to records whose metrics map to our seven metric groups and expand the model axis to the GPT family. We permit built-in web-search here, as leakage from prior literature is negligible for freshly released papers. We collected 52 papers containing 79 previously unseen experimental records and datasets between 2025-09-05 and 2025-09-23. During search, GPT’s web tool may surface the target paper. We match arXiv IDs and verify content to detect any such access, and we exclude those predictions to prevent leakage.
Table 7: Streaming environment prediction results on 79 samples.
Model MAE $\mathbf{r}$
@@ Results
Table 7 summarizes streaming performance for GPT-5 (+Search). On newly released datasets and unseen experiment–dataset combinations, the model achieves accuracy comparable to its performance on Precog (MAE =14.0). The box plot of absolute errors (Figure 5) compares all predictions with the high-confidence subset (>>50% confidence, 23 samples), showing a visibly tighter error distribution for the latter (avg. 9.3). Finally, in a self-referential test that predicted the performance of the task described in this paper before it is indexed by any search engine. GPT-5+Search predicted its own performance at MAE 16.0 versus the realized 14.7 from Table 3. These results indicate that forecasting performance based solely on text-only descriptions could be a practical and achievable approach in real-world applications.
9 Conclusion
We introduced anticipatory evaluation, a setting in which models forecast evaluation metrics on a new task before any experiments are run, using only natural-language descriptions of the proposed tasks and experimental configuration. To investigate this problem, we presented PRECOG, a literature-grounded corpus of 2,290 description-performance pairs curated from 1,519 arXiv papers with careful controls and analysis to mitigate potential data contamination.
Our experiments show that text-only performance forecasting is challenging but feasible: reasoning-enabled LLMs substantially outperform embedding-based baselines, achieving mean absolute error 14.6 overall and 9.9 on high-confidence predictions, with stable performance across post-knowledge-cutoff and streaming evaluations. In a human comparison study, expert forecasters exhibit significantly higher error, highlighting both the intrinsic difficulty of the task and the potential value of LLM-based predictors.
While anticipatory evaluation is not intended to replace empirical benchmarking, our results suggest it has the potential to serve as a lightweight planning tool to aid researchers in better estimating the difficulty of a new task, which could help to prioritize experiments and guide the design of pilot studies before committing annotation, compute, or engineering effort. We hope PRECOG encourages further work on predictive evaluation methods that support earlier, more informed decision-making in machine learning research.
10 Impact Statement
This work aims to advance machine learning methods for forecasting the outcomes of scientific and engineering ideas before expensive real-world experimentation. If successful, such forecasting could help researchers and organizations allocate resources more efficiently, prioritize promising directions earlier, and reduce the time and cost needed to iterate on new hypotheses.
At the same time, forecasting systems built on LLM reasoning and retrieval may introduce risks. Predictions can be overconfident or systematically biased toward well-studied domains, popular research topics, or datasets that are easier to describe, potentially reinforcing existing inequities in what gets explored. Retrieval-based components can also propagate errors or outdated evidence, and model outputs may be misused as a substitute for careful experimental validation. To mitigate these concerns, future work should emphasize uncertainty estimation, calibration, transparency about evidence sources, and evaluation protocols that stress-test performance under distribution shift and sparse prior evaluations (e.g., for new models with few or no reported results).
Finally, improving the robustness and cost-efficiency of forecasting—through stronger priors, selectively collecting a small number of seed evaluations, and using hybrid text–metadata representations—may make these tools more accessible. We hope that developing these components on top of Precog and the streaming benchmark will contribute to more reliable forecasting systems, while encouraging responsible use in settings where predictions inform high-stakes research decisions.
@@ Appendix A Details of Data Curation
For curation we used GPT-5-mini (OpenAI, 2025). The model extracted schema-aligned descriptions from source papers and handled dataset-source retrieval annotations. These choices were based on a pilot run and seed-result quality.
@@
All records involve GPT-4 (Achiam et al., 2023) or GPT-4o (Hurst et al., 2024), fixing the model/method axis to focus on dataset/task variation. We excluded multimodal domains, correlation-only metrics (not on a 0–100 scale), and overall benchmark aggregates (e.g., MMLU) via heuristic filters.
@@ Appendix C Supplementary Experimental Results
We ran additional experiments on the Precog-2025 dataset using more recent embedding models that are competitive or state-of-the-art on the MTEB leaderboard, including Qwen3-8B-Embed (Zhang et al., 2025) and KaLM-Embedding-Gemma3-12B-2511 (Hu et al., 2025), combined with both KNN and XGBoost regressors. The results are summarized in Table 8. As the table shows, newer and stronger embedding models yield modest improvements in MAE and correlation, but the gains are relatively small.